OpenStudy (anonymous):

what is equal to 2^5 x 2^3 ?

OpenStudy (charlotte123):

Is that - 2^5 * 2^3?

OpenStudy (solomonzelman):

when you multiply numbers you add their exponents. \[x^a \times x^b=x^{a+b}\]

OpenStudy (solomonzelman):

\[2^5 \times 2^3=2^{5 + 3}=2^8\]

OpenStudy (charlotte123):

(2 * 2 * 2 * 2 *2) * (2 * 2 * 2) 32 * 8 ----> 256

OpenStudy (anonymous):

yes , charlotte .

OpenStudy (charlotte123):

@SolomonZelman Your method is much easier than mine LOL - Less work/Easier to figure out too *-*

OpenStudy (charlotte123):

Thanks Griffin!

OpenStudy (anonymous):

welcome !

OpenStudy (solomonzelman):

yeah, it's just a matter of knowing the formulaic approaches to these problems.